# Close-Talk, and Pseudo-Label Based Far-Field, Speech Enhancement

### v1

@@ VII-A Results of Close-Talk Speech Enhancement
Row Systems REAL REAL REAL REAL
0 - Close-talk mixture 2.818 349 123 916 3092.8183491239163092.818\,349\,123\,916\,3092.818 349 123 916 309 2.502 779 209 488 744 42.50277920948874442.502\,779\,209\,488\,744\,42.502 779 209 488 744 4 1.143 109 9961.1431099961.143\,109\,9961.143 109 996 1.490 027 5581.4900275581.490\,027\,5581.490 027 558
1 S CTSEnet 3.263 852 279 278 555 43.26385227927855543.263\,852\,279\,278\,555\,43.263 852 279 278 555 4 3.216 589 413 447 3233.2165894134473233.216\,589\,413\,447\,3233.216 589 413 447 323 1.150 484 899 885 6891.1504848998856891.150\,484\,899\,885\,6891.150 484 899 885 689 1.555 420 617 497 314 31.55542061749731431.555\,420\,617\,497\,314\,31.555 420 617 497 314 3
Row $1$ reports the results of CTSEnet trained via supervised learning on the CHiME-4 simulated mixtures (denoted as S in the “Training data” column). Compared with row 00, the DNSMOS scores are clearly improved (e.g., from $2.50$ to $3.22$ on the real test set). The ASR performance becomes worse, possibly because CTSEnet is trained only on simulated data and could hence introduce some distortion to target speech. Nonetheless, the ASR performance degrades rather slightly (e.g., from $1.49$% to $1.56$% WER on the real test set), meaning that the estimated close-talk speech is still of high quality and could be a good pseudo-label for far-field mixtures.
TABLE III: ctPuLSE vs. Purely-Supervised Models on CHiME-4 Far-Field Mixtures (Single-Channel Input)
@@
0 Mixture - - - - - 7.506 217 692 752 9057.5062176927529057.506\,217\,692\,752\,9057.506 217 692 752 905 7.539 180 193 284 8817.5391801932848817.539\,180\,193\,284\,8817.539 180 193 284 881 1.273 397 988 803 4591.2733979888034591.273\,397\,988\,803\,4591.273 397 988 803 459 0.870 243 137 103 452 40.87024313710345240.870\,243\,137\,103\,452\,40.870 243 137 103 452 4 1.520 743 794 895 107 71.52074379489510771.520\,743\,794\,895\,107\,71.520 743 794 895 107 7 1.392 663 582 045 257 81.39266358204525781.392\,663\,582\,045\,257\,81.392 663 582 045 257 8 5.932 890 8555.9328908555.932\,890\,8555.932 890 855 4.070 946 5684.0709465684.070\,946\,5684.070 946 568 8.288 195 7418.2881957418.288\,195\,7418.288 195 741 4.470 082 6754.4700826754.470\,082\,6754.470 082 675
1a Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ 17.295 841 213 880 16317.29584121388016317.295\,841\,213\,880\,16317.295 841 213 880 163 17.665 333 817 529 917.665333817529917.665\,333\,817\,529\,917.665 333 817 529 9 2.361 246 844 584 2922.3612468445842922.361\,246\,844\,584\,2922.361 246 844 584 292 0.959 758 057 782 008 20.95975805778200820.959\,758\,057\,782\,008\,20.959 758 057 782 008 2 3.327 439 949 427 8173.3274399494278173.327\,439\,949\,427\,8173.327 439 949 427 817 3.207 257 983 124 223.207257983124223.207\,257\,983\,124\,223.207 257 983 124 22 3.488 200 589 970 5013.4882005899705013.488\,200\,589\,970\,5013.488 200 589 970 501 2.182 971 348 501 050 82.18297134850105082.182\,971\,348\,501\,050\,82.182 971 348 501 050 8 7.634 478 894 284 6467.6344788942846467.634\,478\,894\,284\,6467.634 478 894 284 646 5.236 115 652 295 763 55.23611565229576355.236\,115\,652\,295\,763\,55.236 115 652 295 763 5
1b Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ 17.229 124 567 725 44217.22912456772544217.229\,124\,567\,725\,44217.229 124 567 725 442 17.647 178 119 958 08817.64717811995808817.647\,178\,119\,958\,08817.647 178 119 958 088 2.489 149 991 310 0042.4891499913100042.489\,149\,991\,310\,0042.489 149 991 310 004 0.962 018 627 202 278 70.96201862720227870.962\,018\,627\,202\,278\,70.962 018 627 202 278 7 3.317 335 051 347 744 63.31733505134774463.317\,335\,051\,347\,744\,63.317 335 051 347 744 6 3.211 204 003 775 436 43.21120400377543643.211\,204\,003\,775\,436\,43.211 204 003 775 436 4 3.392 330 383 480 8263.3923303834808263.392\,330\,383\,480\,8263.392 330 383 480 826 2.087 097 606 843 912.087097606843912.087\,097\,606\,843\,912.087 097 606 843 91 7.410 347 403 810 2367.4103474038102367.410\,347\,403\,810\,2367.410 347 403 810 236 4.442 057 078 798 6364.4420570787986364.442\,057\,078\,798\,6364.442 057 078 798 636
1c Supervised S TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ 17.264 941 603 848 42817.26494160384842817.264\,941\,603\,848\,42817.264 941 603 848 428 17.626 689 404 382 6417.6266894043826417.626\,689\,404\,382\,6417.626 689 404 382 64 2.432 730 621 280 0922.4327306212800922.432\,730\,621\,280\,0922.432 730 621 280 092 0.963 050 892 206 5430.9630508922065430.963\,050\,892\,206\,5430.963 050 892 206 543 3.304 831 562 414 722 43.30483156241472243.304\,831\,562\,414\,722\,43.304 831 562 414 722 4 3.187 231 883 395 354 33.18723188339535433.187\,231\,883\,395\,354\,33.187 231 883 395 354 3 3.366 519 174 041 2983.3665191740412983.366\,519\,174\,041\,2983.366 519 174 041 298 2.238 283 122 534 0172.2382831225340172.238\,283\,122\,534\,0172.238 283 122 534 017 7.209 562 943 593 5757.2095629435935757.209\,562\,943\,593\,5757.209 562 943 593 575 5.166 051 660 516 6055.1660516605166055.166\,051\,$6 051 6$ 605
2 Supervised S iNeuBe [51] - - - $15.1$ $0.954$
3a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ 17.136 097 766 684 76317.13609776668476317.136\,097\,766\,684\,76317.136 097 766 684 763 17.561 826 082 102 69817.56182608210269817.561\,826\,082\,102\,69817.561 826 082 102 698 2.411 083 760 225 411 62.41108376022541162.411\,083\,760\,225\,411\,62.411 083 760 225 411 6 0.961 621 019 343 370 30.96162101934337030.961\,621\,019\,343\,370\,30.961 621 019 343 370 3 3.356 713 550 361 633 73.35671355036163373.356\,713\,550\,361\,633\,73.356 713 550 361 633 7 3.227 691 624 793 467 73.22769162479346773.227\,691\,624\,793\,467\,73.227 691 624 793 467 7 3.160 029 498 525 073 63.16002949852507363.160\,029\,498\,525\,073\,63.160 029 498 525 073 6 1.998 598 768 391 1651.9985987683911651.998\,598\,768\,391\,1651.998 598 768 391 165 6.840 679 865 521 1066.8406798655211066.840\,679\,865\,521\,1066.840 679 865 521 106 3.227 614 554 626 5593.2276145546265593.227\,614\,554\,626\,5593.227 614 554 626 559
3b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 16.710 767 206 369 02416.71076720636902416.710\,767\,206\,369\,02416.710 767 206 369 024 17.254 252 569 589 13817.25425256958913817.254\,252\,569\,589\,13817.254 252 569 589 138 2.412 104 403 972 6262.4121044039726262.412\,104\,403\,972\,6262.412 104 403 972 626 0.960 142 528 725 129 80.96014252872512980.960\,142\,528\,725\,129\,80.960 142 528 725 129 8 3.322 187 319 009 8473.3221873190098473.322\,187\,319\,009\,8473.322 187 319 009 847 3.162 385 252 840 246 53.16238525284024653.162\,385\,252\,840\,246\,53.162 385 252 840 246 5 3.558 259 587 020 6493.5582595870206493.558\,259\,587\,020\,6493.558 259 587 020 649 2.083 410 155 241 712 32.08341015524171232.083\,410\,155\,241\,712\,32.083 410 155 241 712 3 6.854 688 083 675 7576.8546880836757576.854\,688\,083\,675\,7576.854 688 083 675 757 3.563 921 715 166 5193.5639217151665193.563\,921\,715\,166\,5193.563 921 715 166 519
3c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}$ $\mathcal{L}_{X,q}^{\text{real}}$ 16.834 305 574 496 58616.83430557449658616.834\,305\,574\,496\,58616.834 305 574 496 586 17.251 441 973 969 0717.2514419739690717.251\,441\,973\,969\,0717.251 441 973 969 07 2.390 075 166 478 301 82.39007516647830182.390\,075\,166\,478\,301\,82.390 075 166 478 301 8 0.958 571 250 512 548 70.95857125051254870.958\,571\,250\,512\,548\,70.958 571 250 512 548 7 3.347 124 211 800 380 83.34712421180038083.347\,124\,211\,800\,380\,83.347 124 211 800 380 8 3.197 608 615 042 869 33.19760861504286933.197\,608\,615\,042\,869\,33.197 608 615 042 869 3 3.716 814 159 292 0353.7168141592920353.716\,814\,159\,292\,0353.716 814 159 292 035 2.072 347 800 435 119 32.07234780043511932.072\,347\,800\,435\,119\,32.072 347 800 435 119 3 7.639 148 300 336 1977.6391483003361977.639\,148\,300\,336\,1977.639 148 300 336 197 3.778 784 623 289 2713.7787846232892713.778\,784\,623\,289\,2713.778 784 623 289 271
4a ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $256$ 16.712 464 920 499 1116.7124649204991116.712\,464\,920\,499\,1116.712 464 920 499 11 17.167 910 347 776 17517.16791034777617517.167\,910\,347\,776\,17517.167 910 347 776 175 2.293 470 420 349 7282.2934704203497282.293\,470\,420\,349\,7282.293 470 420 349 728 0.957 785 015 609 969 50.95778501560996950.957\,785\,015\,609\,969\,50.957 785 015 609 969 5 3.332 327 093 435 673.332327093435673.332\,327\,093\,435\,673.332 327 093 435 67 3.190 962 144 995 6343.1909621449956343.190\,962\,144\,995\,6343.190 962 144 995 634 3.211 651 917 404 133.211651917404133.211\,651\,917\,404\,133.211 651 917 404 13 1.980 161 510 380 176 41.98016151038017641.980\,161\,510\,380\,176\,41.980 161 510 380 176 4 7.064 811 355 995 5187.0648113559955187.064\,811\,355\,995\,5187.064 811 355 995 518 3.409 780 933 252 373.409780933252373.409\,780\,933\,252\,373.409 780 933 252 37
4b ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $128$ 16.523 509 743 087 2216.5235097430872216.523\,509\,743\,087\,2216.523 509 743 087 22 17.081 067 725 176 98617.08106772517698617.081\,067\,725\,176\,98617.081 067 725 176 986 2.371 422 376 506 256 32.37142237650625632.371\,422\,376\,506\,256\,32.371 422 376 506 256 3 0.959 743 330 086 607 50.95974333008660750.959\,743\,330\,086\,607\,50.959 743 330 086 607 5 3.354 864 545 712 0263.3548645457120263.354\,864\,545\,712\,0263.354 864 545 712 026 3.224 684 022 093 959 43.22468402209395943.224\,684\,022\,093\,959\,43.224 684 022 093 959 4 3.407 079 646 017 699 43.40707964601769943.407\,079\,646\,017\,699\,43.407 079 646 017 699 4 1.876 912 865 518 641.876912865518641.876\,912\,865\,518\,641.876 912 865 518 64 7.083 488 980 201 7197.0834889802017197.083\,488\,980\,201\,7197.083 488 980 201 719 3.199 588 957 914 8963.1995889579148963.199\,588\,957\,914\,8963.199 588 957 914 896
4c ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $64$ 17.016 686 683 125 56717.01668668312556717.016\,686\,683\,125\,56717.016 686 683 125 567 17.537 237 975 786 34317.53723797578634317.537\,237\,975\,786\,34317.537 237 975 786 343 2.402 777 045 513 644 32.40277704551364432.402\,777\,045\,513\,644\,32.402 777 045 513 644 3 0.961 299 180 774 926 30.96129918077492630.961\,299\,180\,774\,926\,30.961 299 180 774 926 3 3.380 592 781 113 3393.3805927811133393.380\,592\,781\,113\,3393.380 592 781 113 339 3.272 828 277 426 903 73.27282827742690373.272\,828\,277\,426\,903\,73.272 828 277 426 903 7 3.289 085 545 722 7143.2890855457227143.289\,085\,545\,722\,7143.289 085 545 722 714 1.869 537 962 314 244 51.86953796231424451.869\,537\,962\,314\,244\,51.869 537 962 314 244 5 6.966 753 828 912 9626.9667538289129626.966\,753\,828\,912\,9626.966 753 828 912 962 3.134 195 898 921 0153.1341958989210153.134\,195\,898\,921\,0153.134 195 898 921 015
4d ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $32$ 16.478 895 149 899 2516.4788951498992516.478\,895\,149\,899\,2516.478 895 149 899 25 17.141 230 481 170 06617.14123048117006617.141\,230\,481\,170\,06617.141 230 481 170 066 2.360 100 678 873 7852.3601006788737852.360\,100\,678\,873\,7852.360 100 678 873 785 0.961 100 110 551 726 20.96110011055172620.961\,100\,110\,551\,726\,20.961 100 110 551 726 2 3.345 090 992 793 603 23.34509099279360323.345\,090\,992\,793\,603\,23.345 090 992 793 603 2 3.224 999 633 733 2563.2249996337332563.224\,999\,633\,733\,2563.224 999 633 733 256 3.384 955 752 212 389 53.38495575221238953.384\,955\,752\,212\,389\,53.384 955 752 212 389 5 1.987 536 413 584 571 61.98753641358457161.987\,536\,413\,584\,571\,61.987 536 413 584 571 6 6.957 415 016 809 8626.9574150168098626.957\,415\,016\,809\,8626.957 415 016 809 862 3.545 237 984 025 413.545237984025413.545\,237\,984\,025\,413.545 237 984 025 41
4e ctPuLSE S+R TF-GridNet $\mathcal{L}_{X,q}^{\text{simu}}+\mathcal{L}_{V,q}^{\text{simu}}+\mathcal{L}_{Y,q}$ $\mathcal{L}_{X,q}^{\text{real}}+\mathcal{L}_{Y,q}$ $16$ 16.078 147 192 231 63316.07814719223163316.078\,147\,192\,231\,63316.078 147 192 231 633 16.928 411 923 202 48216.92841192320248216.928\,411\,923\,202\,48216.928 411 923 202 482 2.364 959 328 012 0332.3649593280120332.364\,959\,328\,012\,0332.364 959 328 012 033 0.959 443 197 977 285 30.95944319797728530.959\,443\,197\,977\,285\,30.959 443 197 977 285 3 3.361 895 135 599 697 33.36189513559969733.361\,895\,135\,599\,697\,33.361 895 135 599 697 3 3.244 502 963 142 3823.2445029631423823.244\,502\,963\,142\,3823.244 502 963 142 382 3.436 578 171 091 4463.4365781710914463.436\,578\,171\,091\,4463.436 578 171 091 446 1.965 411 703 971 385 21.96541170397138521.965\,411\,703\,971\,385\,21.965 411 703 971 385 2 7.228 240 567 799 7767.2282405677997767.228\,240\,567\,799\,7767.228 240 567 799 776 3.040 777 243 215 473.040777243215473.040\,777\,243\,215\,473.040 777 243 215 47
TABLE IV: ctPuLSE vs. Purely-Supervised Models on CHiME-4 Far-Field Mixtures (Six-Channel Input)

### v2

@@ VII.4.1 With or Without Pre-training on Simulated Mixtures
origin=c]0Row Systems REAL REAL SIMU REAL SIMU REAL
0 Mixture - 1.520 743 794 895 107 71.520\,743\,794\,895\,107\,7 1.392 663 582 045 257 81.392\,663\,582\,045\,257\,8 5.932 890 8555.932\,890\,855 4.070 946 5684.070\,946\,568 8.288 195 7418.288\,195\,741 4.470 082 6754.470\,082\,675
1a of Table 4 Supervised S 2.120 078 131 527 103 32.120\,078\,131\,527\,103\,3 1.776 820 183 864 4611.776\,820\,183\,864\,461 0.862 831 858 407 079 60.862\,831\,858\,407\,079\,6 23.164 570 965 006 08423.164\,570\,965\,006\,084 1.340 119 536 794 919 71.340\,119\,536\,794\,919\,7 46.713 998 785 557 47646.713\,998\,785\,557\,476
4c of Table 4 ctPuLSE S+R 3.385 378 211 815 196 33.385\,378\,211\,815\,196\,3 3.289 810 700 044 8533.289\,810\,700\,044\,853 0.859 144 542 772 861 40.859\,144\,542\,772\,861\,4 1.253 733 544 747 225 21.253\,733\,544\,747\,225\,2 1.368 135 973 104 221 21.368\,135\,973\,104\,221\,2 1.700 219 533 840 9081.700\,219\,533\,840\,908
@@
Table 6: Results of ctPuLSE trained with different batch sizes on CHiME-4 far-field mixtures (six-channel input)
@@
Batch size Val. Test Val. Test
origin=c]0Row Systems REAL REAL SIMU REAL SIMU REAL
0 Mixture - 1.520 743 794 895 107 71.520\,743\,794\,895\,107\,7 1.392 663 582 045 257 81.392\,663\,582\,045\,257\,8 5.932 890 8555.932\,890\,855 4.070 946 5684.070\,946\,568 8.288 195 7418.288\,195\,741 4.470 082 6754.470\,082\,675
4c of Table 4 ctPuLSE 1 3.385 378 211 815 196 33.385\,378\,211\,815\,196\,3 3.289 810 700 044 8533.289\,810\,700\,044\,853 0.859 144 542 772 861 40.859\,144\,542\,772\,861\,4 1.253 733 544 747 225 21.253\,733\,544\,747\,225\,2 1.368 135 973 104 221 21.368\,135\,973\,104\,221\,2 1.700 219 533 840 9081.700\,219\,533\,840\,908
6a ctPuLSE 2 3.386 070 151 469 160 53.386\,070\,151\,469\,160\,5 3.287 134 559 354 1473.287\,134\,559\,354\,147 0.833 333 333 333 333 30.833\,333\,333\,333\,333\,3 1.283 233 157 564 8071.283\,233\,157\,564\,807 1.368 135 973 104 221 21.368\,135\,973\,104\,221\,2 1.653 510 205 988 1361.653\,510\,205\,988\,136
6b ctPuLSE 4 3.372 617 047 609 089 53.372\,617\,047\,609\,089\,5 3.246 052 016 567 213 53.246\,052\,016\,567\,213\,5 0.892 330 383 480 8260.892\,330\,383\,480\,826 1.264 795 899 553 818 41.264\,795\,899\,553\,818\,4 1.335 450 130 743 369 41.335\,450\,130\,743\,369\,4 1.709 561 399 411 462 31.709\,561\,399\,411\,462\,3
6c ctPuLSE 6 3.375 136 164 105 228 43.375\,136\,164\,105\,228\,4 3.246 384 706 226 997 73.246\,384\,706\,226\,997\,7 0.881 268 436 578 171 20.881\,268\,436\,578\,171\,2 1.327 482 576 791 179 51.327\,482\,576\,791\,179\,5 1.396 152 409 413 522 71.396\,152\,409\,413\,522\,7 1.728 245 130 552 571 31.728\,245\,130\,552\,571\,3
Table 7: Results of ctPuLSE trained with different weighting term $\alpha$ (for the loss in Eq. (14)) on CHiME-4 far-field mixtures (six-channel input)
@@
Val. Test Val. Test
origin=c]0Row Systems $\alpha$ in Eq. (14) REAL REAL SIMU REAL SIMU REAL
0 Mixture - 1.520 743 794 895 107 71.520\,743\,794\,895\,107\,7 1.392 663 582 045 257 81.392\,663\,582\,045\,257\,8 5.932 890 8555.932\,890\,855 4.070 946 5684.070\,946\,568 8.288 195 7418.288\,195\,741 4.470 082 6754.470\,082\,675
4c of Table 4 ctPuLSE 5 3.385 378 211 815 196 33.385\,378\,211\,815\,196\,3 3.289 810 700 044 8533.289\,810\,700\,044\,853 0.859 144 542 772 861 40.859\,144\,542\,772\,861\,4 1.253 733 544 747 225 21.253\,733\,544\,747\,225\,2 1.368 135 973 104 221 21.368\,135\,973\,104\,221\,2 1.700 219 533 840 9081.700\,219\,533\,840\,908
7a ctPuLSE 10 3.401 449 803 506 894 33.401\,449\,803\,506\,894\,3 3.296 591 379 797 4643.296\,591\,379\,797\,464 0.859 144 542 772 861 40.859\,144\,542\,772\,861\,4 1.261 108 447 951 620 61.261\,108\,447\,951\,620\,6 1.363 466 567 052 6711.363\,466\,567\,052\,671 1.709 561 399 411 462 31.709\,561\,399\,411\,462\,3
7b ctPuLSE 2.5 3.405 142 005 645 373.405\,142\,005\,645\,37 3.302 443 070 999 6963.302\,443\,070\,999\,696 0.870 206 489 675 516 20.870\,206\,489\,675\,516\,2 1.283 233 157 564 8071.283\,233\,157\,564\,807 1.307 433 694 434 0681.307\,433\,694\,434\,068 1.606 800 878 135 363 61.606\,800\,878\,135\,363\,6
7c ctPuLSE 1.0 3.397 188 011 452 2713.397\,188\,011\,452\,271 3.292 664 756 077 5913.292\,664\,756\,077\,591 0.866 519 174 041 2980.866\,519\,174\,041\,298 1.253 733 544 747 225 21.253\,733\,544\,747\,225\,2 1.321 441 912 588 718 81.321\,441\,912\,588\,718\,8 1.728 245 130 552 571 31.728\,245\,130\,552\,571\,3
7d ctPuLSE 0.5 3.390 984 208 319 8013.390\,984\,208\,319\,801 3.289 141 732 859 152 33.289\,141\,732\,859\,152\,3 0.870 206 489 675 516 20.870\,206\,489\,675\,516\,2 1.235 296 286 736 236 61.235\,296\,286\,736\,236\,6 1.382 144 191 258 871 81.382\,144\,191\,258\,871\,8 1.690 877 668 270 353 61.690\,877\,668\,270\,353\,6
7e ctPuLSE 0.25 3.388 038 322 580 9723.388\,038\,322\,580\,972 3.280 015 652 484 611 43.280\,015\,652\,484\,611\,4 0.888 643 067 846 607 70.888\,643\,067\,846\,607\,7 1.268 483 351 156 0161.268\,483\,351\,156\,016 1.442 846 469 929 025 11.442\,846\,469\,929\,025\,1 1.714 232 332 196 739 81.714\,232\,332\,196\,739\,8
7f ctPuLSE 0.1 3.372 480 904 704 950 43.372\,480\,904\,704\,950\,4 3.266 704 502 908 7843.266\,704\,502\,908\,784 0.984 513 274 336 283 30.984\,513\,274\,336\,283\,3 1.360 669 641 210 959 11.360\,669\,641\,210\,959\,1 1.503 548 748 599 178 21.503\,548\,748\,599\,178\,2 1.807 650 987 902 2841.807\,650\,987\,902\,284
Table 8: Comparison of ctPuLSE with other approaches on CHiME-4 far-field mixtures
@@
Joint training? #input mics Val. Test Val. Test
origin=c]0Row Systems Frontend REAL REAL SIMU REAL SIMU REAL
0 Mixture - - 1 1.520 743 794 895 107 71.520\,743\,794\,895\,107\,7 1.392 663 582 045 257 81.392\,663\,582\,045\,257\,8 5.932 890 8555.932\,890\,855 4.070 946 5684.070\,946\,568 8.288 195 7418.288\,195\,741 4.470 082 6754.470\,082\,675
1a IRIS [5] Conv-TasNet 1 - - $5.96$ $4.37$ $13.52$ $12.11$
1b IRIS [5] Conv-TasNet 1 - - $3.16$ $2.03$ $6.12$ $3.92$
2 SuperM2M [36] TF-GridNet 1 3.163 384 205 610 3933.163\,384\,205\,610\,393 3.033 267 061 570 5773.033\,267\,061\,570\,577 3.392 330 3833.392\,330\,383 1.836 350 8971.836\,350\,897 6.569 854 3146.569\,854\,314 3.036 106 313.036\,106\,31
3 ctPuLSE (44c of Table 3) TF-GridNet 1 3.380 592 781 113 3393.380\,592\,781\,113\,339 3.272 828 277 426 903 73.272\,828\,277\,426\,903\,7 3.289 085 545 722 7143.289\,085\,545\,722\,714 1.869 537 962 314 244 51.869\,537\,962\,314\,244\,5 6.966 753 828 912 9626.966\,753\,828\,912\,962 3.134 195 898 921 0153.134\,195\,898\,921\,015
4a MultiIRIS [22] Neural WPD 2 - - $2.28$ $2.06$ $2.30$ $3.63$
4b MultiIRIS [22] Neural WPD 2 - - $2.04$ $1.66$ $2.04$ $2.65$
5 SuperM2M [36] TF-GridNet 2 3.006 918 163 792 940 33.006\,918\,163\,792\,940\,3 2.903 423 564 541 1312.903\,423\,564\,541\,131 1.497 050 1471.497\,050\,147 1.401 231 6081.401\,231\,608 2.082 555 0982.082\,555\,098 1.943 108 0381.943\,108\,038
6 ctPuLSE (44c of Table 4) TF-GridNet 2 3.404 649 122 969 8913.404\,649\,122\,969\,891 3.312 104 813 690 641 73.312\,104\,813\,690\,641\,7 1.574 483 775 811 209 51.574\,483\,775\,811\,209\,5 1.452 855 931 265 902 21.452\,855\,931\,265\,902\,2 2.325 364 213 672 0212.325\,364\,213\,672\,021 2.083 236 022 233 642.083\,236\,022\,233\,64
7a MultiIRIS [22] Neural WPD 6 - - $1.19$ $1.32$ $1.29$ $1.85$
7b MultiIRIS [22] Neural WPD 6 - - $1.22$ $1.33$ $1.24$ $1.77$
8 SuperM2M [36] TF-GridNet 6 2.839 865 817 670 8372.839\,865\,817\,670\,837 2.749 583 063 702 629 72.749\,583\,063\,702\,629\,7 0.829 646 0170.829\,646\,017 1.257 420 9961.257\,420\,996 1.368 135 9731.368\,135\,973 1.606 800 8781.606\,800\,878
9 ctPuLSE (44c of Table 4) TF-GridNet 6 3.385 378 211 815 196 33.385\,378\,211\,815\,196\,3 3.289 810 700 044 8533.289\,810\,700\,044\,853 0.859 144 542 772 861 40.859\,144\,542\,772\,861\,4 1.253 733 544 747 225 21.253\,733\,544\,747\,225\,2 1.368 135 973 104 221 21.368\,135\,973\,104\,221\,2 1.700 219 533 840 9081.700\,219\,533\,840\,908
Note #1\#1: The best scores are highlighted in bold in each of the $1$-, 22- and $6$-channel tasks separately.
Note #2\#2: SuperM2M and ctPuLSE use exactly the same TF-GridNet architecture.
Table 9: Effects of speaker reinforcement